OpenStudy (anonymous):

i have a problem this integral. i keep making a mistake but i can't see it ∫−sin2t+sint−sin3tdt

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2} -\sin^2t+sint-\sin^3t dt\]

OpenStudy (psymon):

Oh, those arepowers, gotcha.

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2} -\sin^2t dt +\int\limits_{0}^{\pi/2} sint dt - \int\limits_{0}^{\pi/2} \sin^3t dt\]

OpenStudy (anonymous):

for the first integral i got \[\frac{ -\pi }{ 4 }\]

OpenStudy (psymon):

Im working the whole problem at a time, so I havent gotten to that. Ill work just that integral

OpenStudy (anonymous):

for the second one -1 and for the third one \[\frac{ 2 }{ 3 }\]

OpenStudy (psymon):

The -pi/4 looks right, but I usually don't apply the limits individually. Lemme finish the problem and see what I get :3

OpenStudy (anonymous):

the problem is that in the end a get both fractions with minus. and in the book only the one with pi has minus

OpenStudy (psymon):

Hmm, I got the same thing first go around.

OpenStudy (psymon):

But yes, it appears we did make a sign mistake, I checked.

OpenStudy (psymon):

Found where I did the stupid sign mistake.

OpenStudy (anonymous):

where?

OpenStudy (psymon):

On the sin^3 part. \[\int\limits_{}^{}-\sin ^{3}xdx\] \[-\int\limits_{}^{}(1-\cos ^{2}x)(sinx)dx\] \[-(\int\limits_{}^{}sinxdx-\int\limits_{}^{}sinxcos ^{2}xdx)\] u = cosx, du = -sinxdx, dx = -du/sinx \[-(\int\limits_{}^{}sinxdx+\int\limits_{}^{}u ^{2}du)\] \[-(-cosx+\frac{ \cos ^{3}x }{ 3 })\] Now of course the cosx we know cancels partwat through, so we end up with \[-\frac{ \cos ^{3}x }{ 3 }\]from 0 to pi/2 \[-0-(-\frac{ 1 }{ 3 })=\frac{ 1 }{ 3 }\] So it does end up being positive.

OpenStudy (anonymous):

thanks

OpenStudy (psymon):

Np :3